If 25 g of C3H8 reacts with excess Oxygen, how many moles of water will be formed?

1
Expert's answer
2022-03-08T14:49:02-0500

C3H8 + 5O2 = 3CO2 + 4H2O;

M(C3H8) = 44 g/mole;

n(C3H8) = m/M = 25/44 = 0.57 moles;

According to the chemical reaction: n(H2O) = 4 * n(C3H8) = 4 * 0.57 = 2.28 moles.

Answer: 2.28 moles.